TIME RESOLVED ICLAS SPECTROSCOPY of HO_{2} RADICAL

\underline{V.A. LOZOVSKY}, V.P. BULATOV, F.N. DZEGILENKO, O.M. SARKISOV, Department of Kinetics and Katalysis, N.N. Semenov Institute of Chemical Physics, Kosygina st. 4, 117977 Moscow, Russia; Yu.V. MATYAGIN, E.A. SVIRIDENKOV, Levedev Physical Institute, Leninsky prospekt 53, 117924 Moscow Russia.

HO_{2} radical absorption spectrum was recorded in the 1.22-1.27 mkm range with the spectral resolution 0.13 cm^{−1} using intracavity laser absorption spectrometer based on F_{2}^{−} color centers LiF laser. The rotationnaly resolved vibronic bands: (0,0,0)-(0,0,0), (0,2,0)-(0,1,0) and (0,0,2)-(0,0,1) of the A′ − A′′ electronic transition were analyzed. The mechanism of the forbiden rotational subbands appearence will be discussed.